OpenStudy (anonymous):

simplify the formula: y=sinx^4-cox^4

OpenStudy (anonymous):

\[a^4-b^4=(a^2+b^2)(a^2-b^2)\] is a start

OpenStudy (anonymous):

then since \(\sin^2(x)+\cos^2(x)=1\) you will get \[\sin^2(x)-\cos^2(x)\]

OpenStudy (anonymous):

and doesnt sin2(x)−cos2(x) equal cos2theta?

OpenStudy (anonymous):

sin^2 x - cos ^2 x = -cos 2x

OpenStudy (anonymous):

thank you!